OpenStudy (anonymous):

2x^2+5x-3=0 solutions for x?

OpenStudy (p0sitr0n):

x = -3 x =0.5

OpenStudy (p0sitr0n):

quadratic formula and all that stuff

OpenStudy (anonymous):

the anwsers are x=-3 , x=2 x=5 x=1/2.....how do u get those two numbers when it can only be one number.

OpenStudy (p0sitr0n):

you can chose one of the 2 roots i would go for -3, as its an integer

OpenStudy (anonymous):

okay then what would 4x^2-16=o be? would that be x=2 x=-2 ?

OpenStudy (p0sitr0n):

4x^2 - 16 =0 4x^2 = 16 x^2 = 4 x1 = 2 x2 = -2 yes, you are right

OpenStudy (p0sitr0n):